OpenStudy (tennisthegreat123):

The permiter of a rectangle is 68 feet. Find the dimensions of the rectangle if the ratio of the length to width is 9 : 8, Which of the following would be the best equation to use to solve the problem? A. 2x + 2x =68 B. (9+8) = 68x C. 9x+8x=68 D. 2(9x) + 2(8x) = 68

563blackghost (563blackghost):

The formula to finding the perimeter is... \(\Huge{2l +2w=P}\) What would be your answer to this question?

OpenStudy (tennisthegreat123):

2(9) + 2(8) = 68?

563blackghost (563blackghost):

Close. Since we only know the ratio to each other we dont really know its actual length and width so instead it would be \(\large{9x}\) and \(\large{8x}\). So what would be the answer?

OpenStudy (tennisthegreat123):

9x + 8x = 68?

563blackghost (563blackghost):

It would actually be the format you first had put it in. \(\Huge{2(9x)+2(8x)=68}\)

563blackghost (563blackghost):

So our answer would be.....?

OpenStudy (tennisthegreat123):

so D is the answer?

563blackghost (563blackghost):

Correct ^^

OpenStudy (tennisthegreat123):

I have a question that correlates to this one @563blackghost

563blackghost (563blackghost):

Just post it up ^^